Chapter 16: We are given 3 charges Q1 Q2-+(1/9)x10°C and Q3-4 С Name (Print) Q3 2 m 1 20 1 m 50° Q2 1) What is the magnitude and direction of the force of Q1 on Q3 and of Q2 on Q3? 2) What is the magnitude and direction of the net force on Q3?
Answer #1

1)Force on Q3 due to Q1 =F1= 9*109*(1/9)*10-9*4/22= 1 N (directed from Q1 to Q3)

Force on Q3 due to Q2=F2=9*109*(1/9)*10-9*4/12= 4 N (directed from Q1 to Q3)

2)horizontal component of F1=1*cos20=0.94 N

vertical component of F1=1*sin20=0.34 N

horizontal component of F2=4*cos50=2.57 N

vertical component of F2=4*sin50 = 3.06 N

total horizontal force = 0.94+2.57=3.51 N

total vertical force = 0.34+3.06=3.4 N

Resultant force = sqrt(3.512+3.42)=4.89 N